OpenStudy (anonymous):

what is th standard form that represents a line with the slope -1 and goes through the point (-3,5) possible answers ...x+y=2 ...-x+y=5 ...y=-x+5 ...y=-x+2

OpenStudy (anonymous):

x+y=2. You use point slope form, y-y1 = m(x-x1). Now, you just substitute, put it into Ax+By=C, and, VAMOOSE!!!!! You're good to go